Skip to main content

AAVSOnet Status, 2019-05-25

1 post / 0 new
HQA
HQA's picture
AAVSOnet Status, 2019-05-25

This has not been a particularly good period for the network!  Listed below is the current status.

BSM_Berry.  Operational.

BSM_Hamren.  Temporarily located at Gary Walker's house, nearly all of the upgraded hardware is configured.  We're waiting for a period when Gary is home and the sky permits to do final polar alignment, and then we'll start operations.  Unfortunately, the weather has been horrible in the NE this spring.

BSM_NH2.  Operational.  The filter wheel was replaced (it appeared to be giving intermittent filter positioning errors).  Gary has swapped the linear actuators that control the roof movement with shorter versions to give better torque.

BSM_NM.  Operational.

BSM_S.  Operational.

BSM_TX.  Still commissioning this system, with filter wheel and mount issues.  Hopefully these will be resolved soon.

OC61.  Currently down, awaiting shutter motor repairs.  This will take another week or two.  We are taking the opportunity to upgrade the computer to Windows10 and the software to the latest releases.  A replacement calibration lamp fiber for the eShel spectrograph has been shipped and should arrive in the next week.

SRO.  Currently down, awaiting a new focus motor.

TMO61.  Currently down.  The computer is not working properly, and dome communication appears to have failed.  We hope to resolve this in another week or so.

Arne

 

AAVSO 49 Bay State Rd. Cambridge, MA 02138 aavso@aavso.org 617-354-0484